### v2

# SMSupplementary Material References

Abstract
Uncertainty quantification in deep learning is crucial for safe and reliable decision-making in downstream tasks. Existing methods quantify uncertainty at the last layer or other approximations of the network which may miss some sources of uncertainty in the model. To address this gap, we propose an uncertainty quantification method for large networks based on variation due to regularization. Essentially, predictions that are more (less) sensitive to the regularization of network parameters are less (more, respectively) certain. This principle can be implemented by deterministically tweaking the training loss during the fine-tuning phase and reflects confidence in the output as a function of all layers of the network. We show that regularization variation (RegVar) provides rigorous uncertainty estimates that, in the infinitesimal limit, exactly recover the Laplace approximation in Bayesian deep learning. We demonstrate its success in several deep learning architectures, showing it can scale tractably with the network size while maintaining or improving uncertainty quantification quality. Our experiments across multiple datasets show that RegVar not only identifies uncertain predictions effectively but also provides insights into the stability of learned representations.
1 INTRODUCTION
Figure 1: Summary of the regularization-based approach to uncertainty quantification that underlies RegVar. Step 1: fit a predictive model by MAP; step 2: fit another model using an infinitesimally regularized MAP; step 3: the difference between the predictions of these two models, normalized by the amount of regularization, is exactly the estimate of the predictive variance.
{mdframed}
// training step to update model parameters theta with training sample (x, y)// n: training data size, lam: scalar hyperparameterf = model(x)loss = cross_entropy(f, y) - log_prior(theta) / n + lam * f.abs().mean() / nloss.backward()optimizer.step()...//inference time for input x_testf_pred_var = (model(x_test) - f_pred_mean).abs() / lam
Figure 2: Example of fine-tuning a model using regularization variation (RegVar). The code required to extend existing fine-tuning to implement RegVar is given in blue. Any auto-differentiation package is sufficient, here we use PyTorch (Paszke et al., 2017).
Bayesian neural networks, where a prior is placed on model weights, offer an opportunity to understand uncertainty and improve model robustness (Wilson, 2020; Papamarkou et al., 2024). However, exact computation of the model-weight posterior and predictive distributions is prohibitive given size and complexity. It is instead appealing to approximate it using the maximum a posteriori (MAP) solution and the curvature thereat. However, even this curvature, characterized by a Hessian, can be challenging to compute.
The Laplace approximation of the posterior distribution uses the (negative) Hessian of the log joint distribution at the MAP (Mackay, 1992). For high dimensional $\theta$ – common in deep neural networks – these steps are a prohibitive computational bottleneck even for an approximate posterior. One remedy is further approximation of the Hessian, but this may ignore important directions of curvature.
In this paper we propose using the principle of variation due to regularization (RegVar), which directly approximates the uncertainty of the predicted mean in the Laplace approximation without explicitly computing and inverting the Hessian (while still assuming it exists). Instead, one additional point estimate is required, the prediction-regularized MAP, derived from the network objective with a small amount of regularization added. Under the assumptions that the Laplace approximation is used, we find that the rescaled difference in network outputs given by the MAP and the prediction-regularized MAP recovers the Laplace variance.
Our contributions in this paper are as follows:
  • We develop a set of novel methods around regularization variation (RegVar) and formalize it as the predictive variance of linearized Laplace.
  • In RegVar, variance is derived from the change in the MAP prediction when adjusting the regularizer. The potential appeal of RegVar is the simplicity of the method (see Figure 1) and implementation (see Figure 2 for an example).
  • We show that RegVar scales up to large networks and compare it to several other popular Bayesian deep learning approaches.
The rest of the paper is organized as follows. In Section 2, we discuss related work. In Section 3, we set up the key aspects of the Laplace approximation and connect it to regularization variation in Section 4. Empirical evaluation is explored in Section 5 before discussing conclusions and future work in Section 6.
2 RELATED WORK
The Laplace approximation was first formulated in BDL for small-scale neural networks (MacKay, 1992). Since then, larger architectures have necessitated Hessian approximations, including Gauss-Newton (Foresee and Hagan, 1997) and Kronecker factorization that places a block diagonal structure on the covariate in line with assumed independence between the layers of a network (Martens and Grosse, 2015).
In recent years, with the aforementioned scalable approximations and renewed appreciation of the advantages of Laplace as a post-hoc uncertainty method (Daxberger et al., 2021), a number of new directions were addressed, including relating neural network loss functions to Gaussian process inference (Khan et al., 2019) and exploring the advantages of locally linearized Laplace in out-of-distribution evaluation (Foong et al., 2019). Underfitting in Laplace (Lawrence, 2001) is mitigated in the GGN approximation by switching to the generalized linear model implied by the local linearization (Immer et al., 2021). Further scaling with linearized Laplace applies neural tangent kernels (Deng et al., 2022) and variational approximation (Ortega et al., 2023). For large models, Laplace can be applied to a low-rank adaptor Yang et al. (2024).
Beyond Laplace-based methods, a number of approaches have been proposed for scalable Bayesian inference in neural networks. Given the breadth of this topic, we give a non-exhaustive account (see Papamarkou et al., 2024 for further discussion). Variational inference approximates the posterior with a parameterized distribution, enabling efficient optimization via stochastic gradient descent (Graves, 2011; Blundell et al., 2015; Shen et al., 2024). Stochastic gradient Langevin dynamics and related approaches that take samples from the optimization trajectory of a neural network use gradient noise to approximate Bayesian updates without explicit posterior evaluation (Welling and Teh, 2011; Mandt et al., 2016; Maddox et al., 2019). Dropout-based uncertainty estimation interprets dropout training as an approximate Bayesian inference method (Gal and Ghahramani, 2016). Ensembles of independently trained networks use multiple predictions to quantify the uncertainty of black-box models (Lakshminarayanan et al., 2017; Osband and Van Roy, 2015).
Owing to the ubiquity of the Hessian of the objective and its computational demands, approaches to avoiding its direct evaluation are developed in areas of research distinct of Laplace. In particular, Hessian-free optimization uses the conjugate gradient method to iteratively update a matrix-vector product as part of an inner loop in Newton’s method second-order optimization (Martens and Sutskever, 2011; Pearlmutter, 1994). Finally, Kallus and McInerney (2022) develop a frequentist approach establishing that infinitesimal regularization of the total log likelihood yields a variance estimate that is asymptotically equivalent to the delta method. While avoiding full Fisher matrix computation, this approach requires evaluating a perturbed estimator per test point and output dimension, leading to computational costs that scale accordingly. It is also a frequentist perspective, in contrast to the Bayesian approach developed in this paper.
3 BACKGROUND
Consider the maximum a posteriori (MAP) solution of a Bayesian neural network $f_{\theta}$ with parameters $\theta$ given data $\left\{(x_{i},y_{i})\right\}_{i=1}^{n}$,
$\hat{\theta}$ $\in\arg_{\theta}\max\mathcal{L}_{\theta}$ (1)
$\mathrm{where\;}\mathcal{L}_{\theta}$ $=\sum_{i=1}^{n}\log p(y_{i}\;|\;f_{\theta}(x_{i}))+\log p(\theta),$ (2)
for some observation likelihood $p(y\;|\;\cdot)$ and prior $p(\theta)$. This is equivalent to optimizing a loss (e.g., squared loss for Gaussian likelihood or cross-entropy for categorical likelihood) regularized by $\log p(\theta)$ and means that computing the MAP of a Bayesian neural network is akin to fitting a vanilla neural network.
Our starting point is the learning of parameters $\theta$ of a deep neural network $f_{\theta}$ as per Eq. 1 and 2. While the representational power and accuracy of the network may be high, in many applications it is crucial to also quantify the uncertainty of predictions, such as in autonomous vehicles, healthcare, or recommendations (Kendall and Gal, 2017; Leibig et al., 2017).
Bayesian deep learning (BDL) provides a framework for analyzing uncertainty in deep learning (Papamarkou et al., 2024). The key components are the prior distribution $p(\theta)$ and likelihood function $p(y\;|\;\theta,x)$ in a family of models indexed by $\theta$, and an i.i.d. dataset $\mathcal{D}=\left\{(x_{i},y_{i})\right\}_{i=1}^{n}$ assumed to be collected from a model in this family. The posterior distribution $p(\theta\;|\;\mathcal{D})$ may be flexibly used as the density (or mass) for the expectation of any downstream prediction depending on $\theta$, e.g. expectation or variance of predicted mean, as part of a larger simulation involving fitted deep models, or interpreting the variance of weights for different layers of the network. The main evaluations of interest in BDL are the mean $\mathbb{E}_{p(\theta\;|\;\mathcal{D})}[f_{\theta}(x)]$ and (co)variance $\mathrm{Var}_{p(\theta\;|\;\mathcal{D})}[f_{\theta}(x)]$ of the prediction for a query vector (or set of query vectors) $x$. These will be the focus of our study in this paper.
BDL is doubly intractable. First, the normalizing factor in the Bayesian posterior is intractable for all but the most trivial networks as it entails a sum over high-dimensional $\theta$. Many techniques have been developed to approximate the posterior – the most common being MAP estimation, variational inference, and Markov chain Monte Carlo – and these vary in their scalability and applicability to deep neural nets (Blundell et al., 2015; Ritter et al., 2018). A key desideratum for approximate inference that is often overlooked is the extent to which it is compatible with existing deep learning frameworks and implementations, which may be supported by years of model selection, tuning, and vast training budgets. The MAP point estimate is the most practical in this sense, and is the basis of the Laplace approximation. Second, even with an approximate posterior to hand, the posterior predictive $p(y\;|\;\mathcal{D},x)$ depends on $f_{\theta}$, meaning that the integral cannot be further simplified. Monte Carlo estimation using samples from the posterior is used to address this intractability.
Laplace Approximation
Let $q$ be the Laplace approximation of the posterior, defined as the second-order Taylor expansion around $\hat{\theta}$,
$\log q(\theta)=\log p(\hat{\theta}\;|\;\mathcal{D})-\frac{1}{2}(\theta-\hat{\theta})^{\top}P(\theta-\hat{\theta}),$ (3)
where $P:=-\nabla\nabla_{\theta}\log p(\theta\;|\;\mathcal{D})$. The first derivative does not appear in Eq. 3 because $\hat{\theta}$ is defined as a local optimum of the posterior, where it exists.
Noticing that Eq. 3 takes the form of a quadratic in $\theta$, we see that $q(\theta)=\mathcal{N}(\hat{\theta},P^{-1})$. If precision matrix $P$ can be calculated and inverted, then the first intractability of approximating the posterior is addressed. In addition, the normal form of $q$ also provides a closed-form solution to the model evidence, supporting model and hyperparameter selection (MacKay, 1992). We next discuss precision calculation in more detail.
Precision Matrix
Distribution $q(\theta)$ requires the covariance matrix which entails calculating and inverting the precision matrix $P$ decomposed as,
$P$ $=-\nabla\nabla_{\theta}\log p(\theta\mid\mathcal{D})$
$=\underbrace{-\nabla\nabla_{\theta}\log p(\theta)}_{\text{Gaussian prior}\implies\text{scalar matrix}}\quad-\underbrace{\nabla\nabla_{\theta}\log p(\mathcal{D}\mid\theta)}_{\text{Hessian}\;H}.$ (4)
The prior term in Eq. 4 usually takes a simple form, e.g. constant scalar for Gaussian prior, so our focus from this point will be on $H$, the Hessian of the likelihood.
The Hessian of the likelihood term in Eq. 4 may be further decomposed as,
$H$ $=\underbrace{\nabla_{f}\log p(y\mid f_{\theta}(x))}_{\text{residual}}{\nabla\nabla_{\theta}f_{\theta}(x)}+$
$\underbrace{\nabla\nabla_{f}\log p(y\mid f_{\theta}(x))}_{\text{observation precision}}\quad\nabla_{\theta}f_{\theta}(x)\nabla_{\theta}f_{\theta}(x)^{\top}.$ (5)
In practice, the precision calculation of the Laplace approximation drops the first term – the network Hessian – in Eq. 5 due to computational constraints. This is known as generalized Gauss-Newton (GGN) and is proportional to the outer product of the network Jacobian (second term in Eq. 5). When GGN is used in approximate inference, it assumes the network takes the form of a generalized linear model. In Laplace, this is equivalent to a local linearization of the network prediction around $\hat{\theta}$,
$\tilde{f}_{\theta}(x)=f_{\hat{\theta}}(x)+\nabla_{\theta}f_{\theta}(x)\biggr{\rvert}_{\hat{\theta}}^{\top}(\theta-\hat{\theta}),$ (6)
which experimentally results in more accurate predictions than the original network $f_{\hat{\theta}}$ when used in the Monte Carlo average for evaluation (Foong et al., 2019; Immer et al., 2021). However, the Hessian, even after simplifying with GGN, is still unwieldy or prohibitive to compute, store, and invert for large networks with high dimensional $\theta$.
4 REGULARIZATION VARIATION (RegVar)
Against this background, we target the same mathematical object as the Laplace approximation, requiring the Hessian to exist but avoiding the need to calculate or invert it. To start, notice that an immediate consequence of Eq. 6 is,
$\tilde{f}_{\theta}\sim\mathcal{N}(f_{\hat{\theta}}(x),\nabla_{\theta}f_{\hat{\theta}}(x)^{\top}P^{-1}\nabla_{\theta}f_{\hat{\theta}}(x)),$ (7)
where $\nabla_{\theta}f_{\hat{\theta}}:=\nabla_{\theta}f_{\theta}\biggr{\rvert}_{\hat{\theta}}$ is used for more compact notation. This is an instance of the Bayesian delta method (Wasserman, 2006), differing from the classic delta method by the inclusion of the prior term in $P$ in Eq. 4.
The posterior predictive depends on the form of the observation likelihood function. For regression tasks, Gaussian observation noise with fixed standard deviation $\sigma$ is the standard, resulting in posterior predictive equal to Eq. 7 plus additional variance term $\sigma^{2}$. For classification tasks, the posterior predictive is non-analytical, requiring Monte Carlo averaging over posterior samples. Notwithstanding the particular form of the posterior predictive distribution, the epistemic uncertainty of the network predictions – meaning, the error due to lack of knowledge – is represented by Eq. 7 and has considerable value in active learning, experimental design, and exploration-exploitation.
We set up the prediction-regularized joint distribution as equal to the joint distribution (Eq. 2) with an additional term proportional to the prediction at query point $x$,
${\mathcal{L}_{\theta}}^{(f(x),\lambda)}$ $=\sum_{i=1}^{n}\log p(y_{i}\;|\;f_{\theta}(x_{i}))+\log p(\theta)+\lambda f_{\theta}(x),$ (8)
for some constant $\lambda\in\mathbb{R}$. The prediction-regularized MAP is defined as,
$\hat{\theta}^{(f(x),\lambda)}\in\arg_{\theta}\max{\mathcal{L}_{\theta}}^{(f(x),\lambda)}.$ (9)
Theorem 4.1.
The derivative of the prediction under the prediction-regularized MAP w.r.t. $\lambda$ recovers the variance-covariance term in Eq. 7,
$\nabla_{\lambda}f_{\hat{\theta}^{(f(x),\lambda)}}(x)\biggr{\rvert}_{\lambda=0}=\nabla_{\theta}f_{\hat{\theta}}(x)^{\top}P^{-1}\nabla_{\theta}f_{\hat{\theta}}(x),$ (10)
assuming $\hat{\theta}^{(f(x),\lambda)}$ and $P^{-1}$ exist, and that ${\mathcal{L}_{\theta}}^{(f(x),\lambda)}$ is continuously differentiable w.r.t. $\theta$.
Proof.
See Appendix A. ∎
A parallel frequentist approach is known as the implicit delta method (Kallus and McInerney, 2022). The relationship in Eq. 10 can also be understood as an application of the implicit function theorem (IFT) by Cauchy, e.g. see (Lorraine et al., 2020) for a modern treatment of IFT.
Algorithm 1 Regularization Variation (RegVar)
  Input: network $f_{\theta}$, fitted MAP $\hat{\theta}$, training data $\mathcal{D}$ of size $n$, evaluation inputs $\{\hat{x}_{i=1}^{m}\}$, scalar $\lambda$, step size schedule $\gamma(\cdot)$
  Warm start $\hat{\theta}^{(\hat{r},\lambda)}\leftarrow\hat{\theta}$
  Initialize step counter $j=0$
  repeat
     Sample training example $(x_{i},y_{i})\sim\mathcal{D}$
     Follow gradient $\hat{\theta}^{(\hat{r},\lambda)}\leftarrow\hat{\theta}^{(\hat{r},\lambda)}+n\gamma(j)\nabla_{\theta}\mathcal{L}_{i}\;\;\;$ // see Eq. 15
     $j\leftarrow j+1$
  until $\hat{\theta}^{(\hat{r},\lambda)}$ converges
  Return: predictive variance function $\hat{\sigma}_{f}^{2}(x):=\frac{1}{\lambda}|f_{\hat{\theta}^{(\hat{r},\lambda)}}(x)-f_{\hat{\theta}}(x)|$
A corollary of Theorem 4.1 is that detecting the local change in prediction as $\lambda$ is increased (or decreased) from 0 yields the variance term of the linearized Laplace approximation in Eq. 7. The local change, i.e. the LHS of Eq. 10, may be calculated by finite differences for some suitably small $\lambda$,
$\frac{1}{\lambda}(f_{\hat{\theta}^{(f(x),\lambda)}}(x)-f_{\hat{\theta}}(x))\underset{\lambda\to 0}{\rightarrow}\nabla_{\lambda}f_{\hat{\theta}^{(f(x),\lambda)}}(x)\biggr{\rvert}_{\lambda=0}.$ (11)
It is natural to consider auto-differentiation to calculate the LHS of Eq. 10. From the fact that the optimum $\hat{\theta}^{(f(x),\lambda)}$ also changes with $\lambda$ it can be seen that auto-differentiation must deal with complex operations on optima and is unlikely to lead to a computational advantage unless used in tandem with additional work-arounds.
Putting it together, the regularization variation (RegVar) approximation is,
$\tilde{f}_{\theta}\sim\mathcal{N}\left(f_{\hat{\theta}}(x),\frac{1}{\lambda}(f_{\hat{\theta}^{(f(x),\lambda)}}(x)-f_{\hat{\theta}}(x))\right),$ (12)
and differs from linearized Laplace (Eq. 7) by the lack of an explicit Hessian. It is apparent by this comparison that RegVar trades the computational and storage expense of dealing with the Hessian for that of a point-wise approximation for each query $x$. This operation may be amortized into the network, which we discuss next.
4.1 Amortized Regularizer
RegVar targets the same predictive variance-covariance as the full linearized Laplace approximation. However, when evaluating the uncertainty of many predictions, the requirement to optimize to the $f$-objective specific to each test point is a bottleneck. In order to scale RegVar in the number of test points, we develop an amortized regularization approach that requires only a single model to represent the $f$-regularized parameters for a set of evaluation points.
Amortization is formulated as follows. Find the regularization term $\hat{r}$ for a finite set of evaluation inputs ${\bf\hat{x}}=\{\hat{x}_{1:m}\}$ that minimizes the average squared L2 norm error of the corresponding $f$-regularized parameters,
$\hat{r}=\arg_{r}\min\sum_{i=1}^{m}\|\hat{\theta}^{(f(x_{i}),\lambda)}-\hat{\theta}^{(r,\lambda)}\|_{2}^{2}.$ (13)
Eq. 13 facilitates the calculation of Eq. 12 using $\hat{\theta}^{(\hat{r},\lambda)}$ instead of $\hat{\theta}^{(f(x),\lambda)}$ for arbitrary $x$ in the domain of $f$.
Theorem 4.2.
The optimal regularization term defined by Eq. 13 is $\hat{r}=\frac{1}{m}\sum_{i=1}^{m}f(\hat{x}_{i})$ when $\hat{\theta}^{(f(x),\lambda)}$ and $P^{-1}$ exist and ${\mathcal{L}_{\theta}}^{(f(x),\lambda)}$ is continuously differentiable w.r.t. $\theta$.
Proof.
See Appendix A. ∎
Theorem 4.2 implies that embedding the average network output in the training objective minimizes the L2 norm of the $f$-regularized objective. Furthermore, we can absorb a sign term into $\lambda$ that depends on the sign of the output function by defining,
$\tilde{\lambda}=\lambda(-1)^{\mathbb{I}[f_{\theta}(x)<0]},$ (14)
where $\mathbb{I}[c]$ is the indicator function that evaluates to 1 when the condition $c$ is true and to 0 otherwise. This equates to using the absolute values of both the RegVar regularizer and predictive variance. Putting it together leads to the amortized objective,
$\mathcal{L}_{i}$ $=\log p(y_{i}\;|\;f_{\theta}(x_{i}))+\frac{1}{n}\log p(\theta)+\frac{\lambda}{nm}\|f_{\theta}({\bf\hat{x}})\|_{1},$ (15)
where the final term in Eq. 15 is the (rescaled) sum of absolute values of the function applied to the $m$ evaluation inputs. We emphasize that this sum of absolute values accommodates networks with multiple outputs, a crucial requirement for practical use.
A stochastic optimization algorithm for amortized RegVar is given in Algorithm 1. The procedure avoids the need to explicitly calculate the Hessian and instead optimizes one parameter vector $\hat{\theta}^{(\hat{r},\lambda)}$. The general form of the objective optimized in the algorithm is Eq. 15. We next highlight a special case and a variant of Eq. 15, both of which admit simplified forms of RegVar.
Special Case: In-Sample Amortization
When the train and test data distributions are the same, the data sampled to optimize the log joint probability of the model may also be used to regularize it, giving rise to the single-sample stochastic gradient,
$\mathcal{L}_{i}^{\mathrm{(IS)}}$ $=\log p(y_{i}\;|\;f_{\theta}(x_{i}))+\frac{1}{n}\log p(\theta)+\frac{\lambda}{nm}|f_{\theta}(x_{i})|.$ (16)
Alternatively, a data augmentation approach is applicable to in-sample training when the likelihood is Gaussian (see Appendix B).
Variant: Parameter Uncertainty Quantification
The same amortized regularization strategy also enables parameter uncertainty quantification by finding the regularizer $\hat{t}$ that minimizes the average squared L2 norm over the $K$ dimensions of parameters,
$\hat{t}=\arg_{t}\min\sum_{k=1}^{K}\|\hat{\theta}^{(\theta_{k},\lambda)}-\hat{\theta}^{(t,\lambda)}\|_{2}^{2}.$ (17)
Theorem 4.2 applies to Eq. 17 in combination with absolute value regularization to provide the amortized objective,
$\mathcal{L}_{i}^{\mathrm{(PU)}}$ $=\log p(y_{i}\;|\;f_{\theta}(x_{i}))+\frac{1}{n}\log p(\theta)+\frac{\lambda}{nm}\|\theta\|_{1},$ (18)
which corresponds to a set of sparsity-inducing Laplace priors over the weights. Thus, when used in concert with a Laplace log prior $\log p(\theta)$, the regularization term for RegVar is absorbed, with $\lambda$ perturbing the precision parameter. In this case, it is particularly convenient to implement RegVar for parameter uncertainty, as one needs only to adjust the strength of the existing regularizer, as detailed in Appendix B.3.
Table 1: Held-Out Predictive Errors with GPT-3, XL GPT-3, and ViT Architectures
Method GPT-3 (125M Param.) XL GPT-3 (1.3B Param.) ViT (86M Param.)
Wikitext IMDB Wikitext IMDB Places365
Negative Log Likelihood ($\downarrow$ is better) ±$\pm 95\%$ Confidence Interval
Pre-Trained $9.781\pm 0.009$ $10.374\pm 0.012$ $8.228\pm 0.013$ $9.886\pm 0.015$ $5.900\pm 0.002$
MAP $1.383\pm 0.027$ $9.681\pm 0.015$ $1.523\pm 0.030$ $9.206\pm 0.020$ $5.871\pm 0.012$
Ensemble $3.925\pm 0.058$ $22.84\pm 0.005$ $5.885\pm 0.010$
MC Dropout $1.410\pm 0.026$ $9.551\pm 0.013$ $1.495\pm 0.029$ $9.137\pm 0.019$ $5.878\pm 0.006$
Laplace Diag $1.481\pm 0.021$ $9.090\pm 0.011$ $1.273\pm 0.021$ $8.797\pm 0.017$ $5.872\pm 0.026$
Laplace Full $1.277\pm 0.021$ $8.774\pm 0.009$ $1.198\pm 0.021$ ${8.047}\pm 0.013$ $5.869\pm 0.001$
RegVar (this paper) ${1.143}\pm 0.021$ ${8.009}\pm 0.008$ ${1.151}\pm 0.020$ $8.138\pm 0.007$ ${5.853}\pm 0.014$
Expected Calibration Error ($\downarrow$ is better) ±$\pm 95\%$ Confidence Interval
Pre-Trained $0.163\pm 0.001$ $0.239\pm 0.001$ $0.233\pm 0.002$ $0.287\pm 0.002$ $0.001\pm 0.000$
MAP $0.052\pm 0.001$ $0.224\pm 0.002$ $0.060\pm 0.001$ $0.240\pm 0.002$ $0.009\pm 0.001$
Ensemble $0.153\pm 0.002$ $0.747\pm 0.004$ $0.002\pm 0.000$
MC Dropout $0.036\pm 0.000$ $0.176\pm 0.002$ $0.054\pm 0.001$ $0.219\pm 0.002$ $0.005\pm 0.000$
Laplace Diag $0.180\pm 0.002$ $0.170\pm 0.001$ $0.072\pm 0.001$ $0.203\pm 0.001$ $0.001\pm 0.000$
Laplace Full $0.082\pm 0.001$ $0.136\pm 0.001$ ${0.024\pm 0.000}$ $0.126\pm 0.001$ ${0.000\pm 0.000}$
RegVar (this paper) ${0.019\pm 0.000}$ ${0.044\pm 0.001}$ $0.047\pm 0.001$ ${0.024\pm 0.001}$ $0.001\pm 0.000$
Figure 3: Calibration curves for the methods on the IMDB held-out evaluation with the XL GPT-3 network. Ideal calibration is plotted with dashed lines. 95% confidence intervals are shown with horizontal markers.
5 EXPERIMENTS
We investigate the empirical performance of RegVar in quantifying uncertainty for large models. The goal is to validate the theoretical developments from Section 3 and compare to baselines in Bayesian deep learning. We next describe the experimental setup.
Experimental Setup
The experiments use a Small GPT-3 Architecture language model with 125 million parameters and a Vision Transformer (Base) with 86 million parameters. To study how the methods scale, we also experiment with a larger XL GPT-3 Architecture comprising 1.3 billion parameters. For the language models, we load the pre-trained model and fine-tune it to Wikitext (wikitext-103-raw) comprising unprocessed text extracted from Wikipedia (Merity et al., 2016) and IMDB a binary sentiment classification dataset comprising movie reviews split evenly between positive and negative sentiments (Maas et al., 2011). For the vision model, we fine-tune the pre-trained model to Places365 (places365), an image dataset of 365 scenes (Zhou et al., 2017). The fine-tuning language datasets were subsampled to 40k documents split into 80% train, 10% validation, and 10% test, while the vision dataset was subsampled to 4k images split into 50% train, 25% validation, and 25% test.
The language models were fine-tuned over 1 epoch with stochastic gradient descent using adaptive learning rates, [^1] Adam with learning rate $2\times 10^{-5}$ (Kingma and Ba, 2017). weight decay $0.01$, [^2] Equivalent to $\mathcal{N}(0,50I)$ prior on network parameters. and batch size 4 with 4 gradient accumulation steps. These experiments were conducted on an instance configured with one NVIDIA A100 GPU with 40 GB memory, 12 CPU cores, and 143 GB of system memory. The vision model was fine-tuned over 5 epochs with the same settings without the gradient accumulation on an instance configured with 4 NVIDIA A10 GPUs each with 24 GB memory, 96 CPU cores, and 380 GB of system memory.
Figure 4: Contour plot of held-out negative log likelihood ($\downarrow$ is better) for Wikitext validation as function of the $\lambda$ used to train RegVar and variance rescaling used during inference. Lines isometric to the identity line are shown as gray dashes.
We compared a number of uncertainty quantification methods,
  • Maximum a posteriori (MAP) optimizes the log joint distribution w.r.t. the network parameters (see Eq. 1).
  • Ensemble Last-Layer performs MAP estimation with multiple heads at the last layer with a shared network before the last layer (Lakshminarayanan et al., 2017; Osband and Van Roy, 2015). During training, heads are randomly initialized and, for each training example, a head is sampled at random to receive an update. During inference, the mean and variance of the population of all head evaluations is used for uncertainty. Memory constraints prohibit approximations beyond the last layer and prevent applying ensembling to the XL GPT-3 architecture.
  • Monte Carlo Dropout (MC Dropout) uses dropout to generate approximate samples of the network during inference time (Gal and Ghahramani, 2016). Since dropout is applied at every layer, it is the only other method (apart from RegVar) that considers uncertainty beyond the last layer.
  • Last-Layer Laplace with Diagonal Covariance (Laplace Diag.) using the inverse of the diagonal of the GGN approximation of the Hessian of the parameters in the last layer of the network.
  • Last-Layer Laplace with Full Covariance (Laplace Full) using the inverse of the GGN approximation of the Hessian of the parameters in the last layer of the network.
  • Regularization Variation (RegVar) uses Algorithm 1.
Further variants of the Laplace approximation can be brought to bear on large models (Daxberger et al., 2021). Last-layer diagonal/full covariance Laplace are used here because we found them, in line with previous studies, to be representative of the performance of different Laplace approaches.
Hyperparameters
All network weights are regularized by a Gaussian prior with zero mean and low precision (0.02) during fine-tuning. Due to computational constraints, 10 samples are used in the ensemble and MC dropout methods. Dropout rate, prior conditioning matrix for Laplace, and the $\lambda$ weight in RegVar are tuned to maximize likelihood on validation data. For all methods there is a trade-off between accuracy and calibration; to control for the confounding influence of this trade-off on the metrics, we rescale the predicted variance by a scalar hyperparameter that is chosen for each method to maximize likelihood on the validation datasets.
Evaluation
All methods are evaluated by held-out negative log likelihood (NLL) and expected calibration error (ECE). NLL penalizes predictive uncertainties that place more probability on the wrong classes, indicating both accuracy and calibration. ECE measures calibration directly using the weighted average error of binned predictions,
$\text{ECE}=\sum_{m=1}^{M}\frac{|B_{m}|}{n}\left|\text{acc}(B_{m})-\text{conf}(B_{m})\right|,$ (19)
where $M$ is the total number of bins, $B_{m}$ is the set of samples whose predicted confidence scores fall into the $m$-th bin, $n$ is the total number of samples, $\text{acc}(B_{m})$ is the accuracy of the samples in the $m$-th bin, and $\text{conf}(B_{m})$ is the average predicted confidence of the samples in the $m$-th bin. We use $M=10$ throughout. Confidence intervals are calculated using ±$\pm 1.96$ standard errors (Wald confidence intervals).
For every method, the first step to evaluating the predictive distribution over classes is to estimate the logit means $f_{\theta}(x_{i})$ and variances $\hat{\sigma}^{2}_{i}$ (rescaled by the aforementioned hyperparameter). Uncertainty in the logits is then used in the extended probit approximation (Gibbs, 1998) which introduces a logit correction vector $\kappa$ to account for model variance. The probit approximation correction is given by,
$\kappa_{i}=\frac{1}{\sqrt{1+\frac{\pi\hat{\sigma}^{2}_{i}}{8}}}.$ (20)
The adjusted logits are then used to form predicted probabilities over the clases $p(y_{i}\;|\;\mathcal{D},x_{i})\propto\exp\{\kappa_{i}f_{\theta}(x_{i})\}$.
Figure 5: Fine-tune training and inference times for the methods on IMDB dataset. N.B. ensemble was prohibited in the larger model due to memory constraints.
The implementation was done in PyTorch (Paszke et al., 2017) using HuggingFace333huggingface.co to load pre-trained models from EleutherAI/{gpt-neo-125M,gpt-neo-1.3B} repositories (Black et al., 2021) originally trained on a 800GB text dataset (Gao et al., 2020) and google/vit-base-patch16-224-in21k repositories (Wu et al., 2020) originally trained on ImageNet-21k (Deng et al., 2009). The code to run the experiments can be accessed publicly on GitHub. [^4] https://github.com/jamesmcinerney/regvar
Results
A comparison of evaluation metrics for the methods is shown in Table 3. We find that RegVar offers significant advantages over MAP across datasets and improves or maintains uncertainty quantification in comparison to the other methods. While Laplace Full (Last Layer) performs robustly across settings, improvements with RegVar highlight the potential benefits of deep uncertainty, i.e., including the sources of variance from the entire network, which is not tractable to target with explicit Hessians at this scale. While NLL measures a combination of accuracy and calibration, ECE isolates the contribution of calibration and allows us to explore how this breaks down further with calibration curves in Figure 3. Overall calibration is good, as evidenced by the small deviation from the ideal line, however most methods tend to over-represent extreme predictions (at 0/1) while RegVar is able to calibrate the predictions away from the extremes.
We explore the empirical behavior of the $\lambda$ hyperparameter in RegVar by running fine-tuning with a wide range of different values. Figure 4 shows a contour plot of the effect of $\lambda$ (x-axis) and the variance rescaling hyperparameter (y-axis), the latter a hyperparameter used in all methods. The isometry of the contours suggests insensitivity of held-out negative log likelihood to the choice of $\lambda$ during fine-tuning. The two factors appear to cancel each other out and leave a dataset-dependent constant rescaling term (approx. $\frac{1}{100}$ for Wikitext) that does not depend on $\lambda$.
A wall clock time comparison for fine-tuning and inference is shown in Figure 5. While inference performance of RegVar is advantageous to all other methods apart from MAP, obtaining the regularized network in RegVar doubles the MAP fine-tuning time, a cost that is amenable to further optimization (e.g. warm starting from MAP) and parallelization in a similar fashion to ensembling and dropout.
Additional experiments are detailed in Appendix C, exploring further settings including using the RegVar framework for parameter uncertainty.
Limitations
RegVar is inherently a local method centered around the optimum of the joint distribution and the curvature is evaluated at one location (ignoring other modes). Early stopping during fine-tuning or pre-training also means that the chosen network parameters may not be around the mode of the explicit loss. The finite differences required by RegVar are an additional source of error making it important to select $\lambda$ small enough, but not so small that underflow errors arise or, more practically, that the noise in the gradients overwhelm the signal from the regularizer.
6 CONCLUSIONS & FUTURE WORK
In this paper, we explored a new concept in Bayesian deep learning that identifies variation due to regularization as an approximation to the predictive variance. We considered this phenomenon from a linearized Laplace perspective and evaluated it in large models. There is further work to explore in what cases it is necessary to use the uncertainty of the whole network for predictions to help contextualize the value of RegVar.
ACKNOWLEDGEMENTS
We thank the anonymous AISTATS reviewers for their insightful feedback and suggestions.
Appendix A PROOFS
We detail here the proofs of the theorems given in the main text. Table 2 summarizes the notation used.
Theorem 4.1 The derivative of the prediction under the prediction-regularized MAP w.r.t. $\lambda$ recovers the variance-covariance term in Eq. 7,
$\nabla_{\lambda}f_{\hat{\theta}^{(f(x),\lambda)}}(x)\biggr{\rvert}_{\lambda=0}=\nabla_{\theta}f_{\hat{\theta}}(x)^{\top}P^{-1}\nabla_{\theta}f_{\hat{\theta}}(x),$ (21)
assuming $\hat{\theta}^{(f(x),\lambda)}$ and $P^{-1}$ exist, and that ${\mathcal{L}_{\theta}}^{(x,\lambda)}$ is continuously differentiable w.r.t. $\theta$.
Proof.
To improve readability we use $(x,\lambda)$ in this proof to refer to fact that evaluation $x$ is passed to the regularizing term (i.e. always $f(x)$). Consider the derivative w.r.t. $\theta$ of the prediction-regularized objective at the optimum,
$\left.\nabla_{\theta}{\mathcal{L}_{\theta}}^{(x,\lambda)}\right|_{\theta=\hat{\theta}^{(x,\lambda)}}$
$=\nabla_{\theta}\left.\mathcal{L}_{\theta}\right|_{\theta=\hat{\theta}^{(x,\lambda)}}+\lambda\nabla_{\theta}\left.f_{\theta}(x)\right|_{\theta=\hat{\theta}^{(x,\lambda)}}$
$=0,$ (22)
by definition in Eq. 8 and by definition of the optimum. Now, differentiate Eq. 22 w.r.t. $\lambda$, applying the chain rule and noticing that $\hat{\theta}^{(x,\lambda)}$ is a function of $\lambda$,
$\nabla^{2}_{\theta}\left.\mathcal{L}_{\theta}\right|_{\theta=\hat{\theta}^{(x,\lambda)}}\nabla_{\lambda}\hat{\theta}^{(x,\lambda)}+\nabla_{\theta}\left.f_{\theta}(x)\right|_{\theta=\hat{\theta}^{(x,\lambda)}}$
$+\lambda\nabla^{2}_{\theta}\left.f_{\theta}(x)\right|_{\theta=\hat{\theta}^{(x,\lambda)}}\nabla_{\lambda}\hat{\theta}^{(x,\lambda)}=0.$ (23)
Evaluating Eq. 23 at $\lambda=0$ eliminates the term involving the Hessian of the output. Since, by assumption, the inverse Hessian of the log joint exists, rearrange the terms,
$\nabla_{\lambda}\left.\hat{\theta}^{(x,\lambda)}\right|_{\lambda=0}=-\left(\nabla^{2}_{\theta}\left.\mathcal{L}_{\theta}\right|_{\theta=\hat{\theta}^{(x,0)}}\right)^{-1}\nabla_{\theta}\left.f_{\theta}(x)\right|_{\theta=\hat{\theta}^{(x,0)}}$ (24)
Finally, substituting Eq. 24 into the chain rule expansion of $\nabla_{\lambda}\left.f_{\hat{\theta}^{(x,\lambda)}}(x)\right|_{\lambda=0}$ yields,
$\nabla_{\lambda}\left.f_{\hat{\theta}^{(x,\lambda)}}(x)\right|_{\lambda=0}$
$=-\nabla_{\theta}\left.f_{\theta}(x)\right|_{\theta=\hat{\theta}}^{\top}\left(\nabla^{2}_{\theta}\left.\mathcal{L}_{\theta}\right|_{\theta=\hat{\theta}}\right)^{-1}\nabla_{\theta}\left.f_{\theta}(x)\right|_{\theta=\hat{\theta}}.$ (25)
This recovers Eq. 10 by substituting with the definition of $P$ (Eq. 4). ∎
Theorem 4.2 The optimal regularization term defined by Eq. 13 is $\hat{r}=\frac{1}{M}\sum_{m=1}^{M}f(\hat{x}_{m})$ when $\hat{\theta}^{(f(x),\lambda)}$ and $P^{-1}$ exist and ${\mathcal{L}_{\theta}}^{(f(x),\lambda)}$ is continuously differentiable w.r.t. $\theta$.
Proof.
Take the derivative of the objective in Eq. 13 w.r.t. $r$,
$\nabla_{r}\sum_{m=1}^{M}\|\hat{\theta}^{(f(x_{m}),\lambda)}-\hat{\theta}^{(r,\lambda)}\|_{2}^{2}$
$=\nabla_{r}\sum_{m=1}^{M}\|(\hat{\theta}^{(f(x_{m}),\lambda)}-\hat{\theta})-(\hat{\theta}^{(r,\lambda)}-\hat{\theta}))\|_{2}^{2}$
$=\nabla_{r}\sum_{m=1}^{M}\|P^{-1}\nabla_{\theta}f(\hat{x}_{m})-P^{-1}\nabla_{\theta}r\|_{2}^{2}$ (26)
$=\sum_{m=1}^{M}\nabla_{r}\|P^{-1}(\nabla_{\theta}f(\hat{x}_{m})-\nabla_{\theta}r)\|_{2}^{2}$
$=2(P^{-1})^{\top}P^{-1}(\sum_{m=1}^{M}(\nabla_{\theta}f(x_{m})-\nabla_{\theta}r)),$ (27)
using Theorem 4.1 to express the difference in optimal parameters in terms involving the inverse Hessian in Eq. 26. Finally, set the derivative in Eq. 27 to zero and solve for $r$, the unique minimizer of the objective (up to a constant that does not depend on $\theta$). ∎
Table 2: Notation Summary
Symbol Definition
$\theta$ Network parameters.
$\mathcal{L}_{\theta}$ Log joint distribution parameterized by $\theta$.
$\hat{\theta}$ Maximum a posteriori parameters under the joint distribution (Eq. 1).
$f_{\hat{\theta}}$ Network output with parameters $\hat{\theta}$.
$\nabla_{\theta}f_{\hat{\theta}}$ Derivative of network output w.r.t. $\theta$ evaluated at $\theta=\hat{\theta}$.
${\mathcal{L}_{\theta}}^{(t_{\theta},\lambda)}$ Log joint distribution parameterized by $\theta$ and regularized by $\lambda t_{\theta}$, e.g. $t_{\theta}$ could be a network output $f_{\theta}(x)$ (Eq. 8) or fine-tuning regularizer $r$ (Eq. 13).
$\hat{\theta}^{(t_{\theta},\lambda)}$ Parameters that maximize ${\mathcal{L}_{\theta}}^{(t_{\theta},\lambda)}$, e.g., see Eq. 9.
$\nabla_{\lambda}f_{\hat{\theta}^{(t_{\theta},\lambda)}}$ Derivative of network output under parameters $\hat{\theta}^{(t_{\theta},\lambda)}$ w.r.t. regularization weight $\lambda$. Note that the optimal network parameters themselves are a function of $\lambda$, making the overall term a non-linear expression involving $\nabla_{\lambda}f_{\hat{\theta}}$ (see Eq. 21).
Appendix B ALGORITHMS FOR REGULARIZATION VARIATION (RegVar)
In this section we expand on variations of the RegVar strategy and provide relevant algorithms where applicable.
B.1 Pointwise RegVar
The algorithm that follows most directly from Theorem 1 targets the uncertainty calculation for an arbitrary evaluation point $x$ (not necessarily from the training or test set) and is given in Algorithm 2. The computational burden of recalculating the optimal regularized parameters for every different evaluation input means that this approach is not scalable to large models, motivating extensions to the regularization term which we consider next.
Algorithm 2 Pointwise RegVar
  Input: network $f_{\theta}$, fitted MAP $\hat{\theta}$, training data $\mathcal{D}$ of size $n$, evaluation input $x$, scalar $\lambda$, step size schedule $\gamma(\cdot)$
  Warm start $\hat{\theta}^{(x,\lambda)}\leftarrow\hat{\theta}$
  Initialize step counter $j=0$
  repeat
     Sample training example $(x_{i},y_{i})\sim\mathcal{D}$
     Follow stochastic gradient $\hat{\theta}^{(x,\lambda)}\leftarrow\hat{\theta}^{(x,\lambda)}+N\gamma(j)\nabla_{\theta}\tilde{\mathcal{L}}_{\theta}^{(x,\lambda)}(x_{i},y_{i})$
     $j\leftarrow j+1$
  until $\hat{\theta}^{(x,\lambda)}$ converges
  Return: predictive variance $\hat{\sigma}_{f}^{2}(x):=\frac{1}{\lambda}|f_{\hat{\theta}^{(x,\lambda)}}(x)-f_{\hat{\theta}}(x)|$
B.2 Data-Augmentation Objective
When the likelihood is Gaussian, the RegVar regularization term can be absorbed into the likelihood. In particular, completing the square in Eq. 16, by pulling the term $\frac{\lambda\sigma^{2}}{nM}(\frac{\lambda\sigma^{2}}{2nM}-y_{i})$ from a constant that does not depend on $\theta$, yields the objective,
$\mathcal{L}_{i}^{\mathrm{(DA)}}$ $=\log\mathcal{N}\left(y_{i}+\frac{\tilde{\lambda}}{nM}\sigma^{2}\;\bigr{|}\;f_{\theta}(x_{i}),\sigma^{2}\right)+\frac{1}{n}\log p(\theta),$ (28)
for observation variance hyperparameter $\sigma^{2}$. Under the common Gaussian likelihood assumption, Eq. 28 implies that pre-trained RegVar is performed by data augmentation, without requiring any adjustment to the model architecture or training procedure.
Algorithm 3 Amortized RegVar for Parameter Uncertainty
  Input: network $f_{\theta}$, fitted MAP $\hat{\theta}$, training data $\mathcal{D}$ of size $n$, evaluation inputs $\{\hat{x}_{i=1}^{m}\}$, scalar $\lambda$, step size schedule $\gamma(\cdot)$
  Warm start $\hat{\theta}^{(\hat{r},\lambda)}\leftarrow\hat{\theta}$
  Initialize step counter $j=0$
  repeat
     Sample training example $(x_{i},y_{i})\sim\mathcal{D}$
     Follow gradient $\hat{\theta}^{(\hat{r},\lambda)}\leftarrow\hat{\theta}^{(\hat{r},\lambda)}+n\gamma(j)\nabla_{\theta}\mathcal{L}_{i}^{\mathrm{(PU)}}\;\;\;$ //See Eq. 18
     $j\leftarrow j+1$
  until $\hat{\theta}^{(\hat{r},\lambda)}$ converges
  Return: parameter variance $\hat{\sigma}_{\theta}^{2}(x):=\frac{1}{\lambda}|\hat{\theta}^{(\hat{r},\lambda)}-\hat{\theta}|$
B.3 RegVar for Parameter Uncertainty
For completeness, we present Algorithm 3 that corresponds to the parameter uncertainty objective given in Eq. 18 in the main text. We apply this algorithm in Section C.2.
Appendix C FURTHER EXPERIMENTS AND HYPERPARAMETER SELECTION
In this section, we evaluate RegVar and compare against the full Hessian and its approximations.
C.1 Synthetic Data
We explore four data generating distributions designed to test both in-distribution inference and extrapolation to “in-between” uncertainty (Foong et al., 2019). A feedforward neural network is used with 50 hidden units and $\tanh$ activations to ensure that the Hessian exists, and optimized using Adam (Kingma and Ba, 2017) with a learning rate of 0.005. Four methods for quantifying the epistemic uncertainty of the network are compared,
  • Exact Hessian using Eq. 4 and both terms in the likelihood Hessian (Eq. 5).
  • GGN using Eq. 4 with only the second term in likelihood Hessian (Eq. 5).
  • Eigenvector approximation of the GGN for $P^{-1}\approx A\Lambda^{-1}A^{\top}$ using the top $k$ eigenvectors $A$ of $P$ (with corresponding eigenvalues $\Lambda$) where $k=\log_{e}(\mathrm{dim}(\theta))$ rounded to the nearest positive integer. The GGN is chosen because it only makes sense to use the eigenvector approximation when the Hessian is large.
  • RegVar (this paper), using Algorithm 1.
The metrics used to evaluate uncertainty performance are,
  • Prediction interval coverage probability (PICP) assesses calibration by evaluting how well prediction intervals cover the observed values. It is calculated as the proportion of times the held-out observed value $y$ falls within the 95% confidence interval of the predicted observations (closer to 0.95 is better).
  • Continuous Ranked Probability Score (CRPS) is a generalization of the Brier score for continuous outcomes, measuring the difference between the CDF of the predicted observations and the CDF of the observed response variable (smaller is better).
  • Negative log likelihood (NLL) evaluates the negated predicted log likelihood of the held out observations (smaller is better).
(a) Full Hessian
(b) GGN
(c) Eigen. Approx.
(d) HFL (this paper)
Figure 6: Illustration of epistemic uncertainties of the predicted mean for the Sin-Inbetween OOD task.
The synthetic data-generating distributions are,
  • Quadratic-Uniform Draw 32 data points $x\sim\mathcal{N}(0,1)$ and let $y=\frac{1}{10}x^{2}-\frac{1}{2}x+5+\frac{1}{10}\epsilon$, where $\epsilon\sim\mathcal{N}(0,1)$.
  • Quadratic-Inbetween Draw 16 data points $x\sim\mathrm{Uniform}(-2,-\frac{1}{2})$ and 16 data points $x\sim\mathrm{Uniform}(\frac{4}{5},\frac{5}{2})$, using the same response distribution as Quadratic-Uniform.
  • Sin-Uniform Sample 160 data points $x\sim\mathrm{Uniform}(-\frac{3}{2},\frac{23}{20})$ and let $y=-\sin(3x-\frac{3}{10})+\frac{1}{10}\epsilon$, where $\epsilon\sim\mathcal{N}(0,1)$.
  • Sin-Inbetween Fix 160 data points evenly spaced in ranges $[-1.5,-0.7)$ and $[0.35,1.15)$ and let use the same response as Sin-Uniform.
(a) Hyperparameter sweeps for observational variance.
(b) Lambda sweep
Figure 7: Hyperparameter selection for modeling and RegVar algorithm.
Datasets
Quadratic Quadratic-Inbetween Sin Sin-Inbetween
In-Distribution OOD In-Distribution OOD In-Distribution OOD In-Distribution OOD
Full Hessian PICP: 0.9375 PICP: 0.9556 PICP: 0.9688 PICP: 0.8667 PICP: 1.0000 PICP: 0.8625 PICP: 1.0000 PICP: 0.9250
CRPS: 0.003156 CRPS: 0.009017 CRPS: 0.003797 CRPS: 0.003331 CRPS: 0.04666 CRPS: 0.07498 CRPS: 0.04486 CRPS: 0.06022
NLL: -23.924 NLL: -38.027 NLL: -23.290 NLL: -24.538 NLL: -75.764 NLL: 26.004 NLL: -73.178 NLL: 10.029
GGN PICP: 0.9375 PICP: 0.9556 PICP: 0.9688 PICP: 0.8444 PICP: 1.0000 PICP: 0.8250 PICP: 1.0000 PICP: 0.8875
CRPS: 0.003762 CRPS: 0.007771 CRPS: 0.003992 CRPS: 0.0008355 CRPS: 0.04652 CRPS: 0.05493 CRPS: 0.04482 CRPS: 0.04302
NLL: -23.790 NLL: -38.751 NLL: -23.135 NLL: -23.504 NLL: -76.025 NLL: 101.6 NLL: -73.260 NLL: 24.903
Eigen. Approx. PICP: 0.9375 PICP: 0.9556 PICP: 0.9688 PICP: 0.8444 PICP: 1.0000 PICP: 0.8250 PICP: 1.0000 PICP: 0.8875
CRPS: 0.003760 CRPS: 0.007705 CRPS: 0.003566 CRPS: 0.0001922 CRPS: 0.04605 CRPS: 0.05364 CRPS: 0.04454 CRPS: 0.04208
NLL: -23.789 NLL: -38.760 NLL: -23.098 NLL: -22.893 NLL: -77.027 NLL: 109.7 NLL: -73.785 NLL: 27.112
RegVar PICP: 0.9375 PICP: 0.9333 PICP: 0.8750 PICP: 0.8444 PICP: 1.0000 PICP: 0.8375 PICP: 1.0000 PICP: 0.9250
CRPS: 0.002347 CRPS: 0.006976 CRPS: 0.001782 CRPS: 4.735e-05 CRPS: 0.04691 CRPS: 0.06397 CRPS: 0.05016 CRPS: 0.06301
NLL: -23.899 NLL: -38.440 NLL: -22.571 NLL: -23.676 NLL: -75.245 NLL: 45.483 NLL: -63.218 NLL: 22.915
Table 3: Uncertainty quantification performance across datasets and methods. Full Hessian (top line) is intractable at scale and included here for reference.
For both the quadratic and sin datasets, the observational variance hyperparameter was selected through grid search over the set $\{0.005,0.01,0.05,0.1,0.5,1.0\}$ minimizing negative log likelihood for a held-out validation set of equal size to the training dataset. The results are summarized in Figure 7(a), with the optimal hyperparameter chosen indicated by a marker. The prior variance was fixed at 3.0, which was the minimum value to the nearest 0.1 for which the Hessian was sufficiently conditioned across all tasks to allow for inversion.
Figure 8: The uncertainty ranges of a random subset of 30 parameters of the neural network indicate that many parameters include 0 in their intervals. RegVar recovers much of their structure just via additional regularization of the network parameters in the objective function.
Figure 9: Absolute errors in predicted means for all the tasks with networks with added sparsity using the uncertainty quantification methods.
The results of the evaluations are summarized in Table 3. We find that RegVar performs commensurately with other approximations, and indeed the full Hessian, despite taking a significantly different approach. In particular, in several cases it achieves the best out-of-distribution (OOD) performance, likely due to its ability to avoid the GGN assumption that ignores an important part of the curvature around the MAP. To illustrate this point, we plot the predicted mean output for the methods on the Sin-Inbetween OOD task in Figure. 6. There is a noticeable difference to the GGN and eigenvector approximation, implying that in practice the residual of the loss may not be close to zero.
C.2 Parameter Uncertainty Experiment
We also evaluated the parameter regularization approach given in Eq. 18. Figure 8 shows the 1 standard deviation interval about the predicted mean under the various methods. Recall that RegVar for parameter uncertainty requires only to apply a small amount of additional L1 regularization to the parameters. Nonetheless, the intervals indicated often align with those requiring full Hessian inversion. These parameter uncertainties may be used for model introspection. As an example, we considered the task of adding sparsity to a network, setting parameters to 0 when their uncertainty intervals overlap with 0. The mean absolute errors of the resulting predictions are shown in Figure 9. RegVar does competitively, particularly for in-between uncertainty. An outlier result is the absolute error of the full Hessian in the Quadratic-Inbetween in-distribution task. This was due to mis-selecting an entire group of units to be set to 0, greatly impacting the shape of predictions.